OpenStudy (anonymous):

Dose anyone know what to do to solve this? :( If h(x) = −4x − 10, find h(−5). −30 1.25 10 −1

OpenStudy (anonymous):

susbstitute (-5) for x

OpenStudy (anonymous):

So the quation is -4*5-10?

OpenStudy (anonymous):

-4*-5-10

OpenStudy (anonymous):

oh ok so what now i solve that? would it be 10?

OpenStudy (anonymous):

oh ok what about this one? Given the function f(x) = 2(x + 6), find x if f(x) = 22. 56 5 38 17

OpenStudy (anonymous):

hmm... i think that would be just be 2(22+6)

OpenStudy (anonymous):

ok ty and last one :P Given the function f(x) = −2x + 8, find x if f(x) = 14 −20 36 −3 −11

OpenStudy (anonymous):

yeah, same as the others -2(14+8) think of f(x) being y so sorta like the equation y= -2x+8 if that made any sense

OpenStudy (anonymous):

I dont get how to solve that one

OpenStudy (anonymous):

im sorry, i wrote the equation wrong i meant -2(14) +8

OpenStudy (anonymous):

ok thanks so much!